PHOTOIONIZATION AND FRAGMENTATION DYNAMICS OF ClN_{3}

NILS HANSEN AND ALEC M. WODTKE, Department of Chemistry and Biochemistry, University of California Santa Barbara, Santa Barbara, CA 93106; \underline{ANATOLY V. KOMISSAROV} AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, Georgia 30322.

The photoionization and fragmentation dynamics of ClN_{3} have been examined using 203 nm excitation with (2+1) REMPI detection of the N_{2} product. Kinetic energy and angular distributions of N_{2} and NCl^{+} were characterized by velocity map imaging. The N_{2} product was formed with appreciable rotational excitation, with population in levels as high as J=90. Velocity map images for the J=50 fragment showed that the maximum energy released to translation was 1.14 eV. This result indicated that the N_{2} did not come from the expected channel, ClN_{3} +hν ⟶ NCl(a^{1}Δ) + N_{2}, which would produce much more energetic fragments. Velocity maps of N_{2} and NCl^{+} were consistent with the process

ClN_{3} + 2hν ⟶ NCl^{+} + N_{2}

The observed kinetic energy distributions of the N_{2} and NCl^{+} photoproducts are consistent with the formation of vibrationally excited NCl^{+}. The velocity maps of photoelectrons peaked near zero velocity, showing that ClN_{3}^{+} is formed with nearly all excess energy in vibration. Ab initio calculations (CCSD(T)/cc-pVTZ) confirm that ClN_{3}^{+} is unstable with respect to decomposition to NCl^{+} and N_{2}. In combination, the experimental and theoretical results can be used to obtain the thermodynamics of the ClN_{3} +hν ⟶ NCl(a^{1}Δ) + N_{2} reaction. The fact that products correlating with NCl(a) could not be observed suggests that the state of ClN_{3} accessed by 203 nm excitation does not undergo direct dissociation.